OpenStudy (anonymous):

****PLEASE HELP**** Admira is painting a rectangular banner 2 1/4 yards wide on a wall in the cafeteria. The Banner will have a blue background. Admira has enough blue paint to cover 1 1/2 square yards of wall. Find the height of the banner if Admira uses all of the blue paint.

OpenStudy (confluxepic):

\[\huge \color{blue}{Welcome~To~OpenStudy!!!}\]

OpenStudy (paxpolaris):



OpenStudy (paxpolaris):

\[Area= h \times w\]

OpenStudy (paxpolaris):

\[1\frac12=h \times 2 \frac14\]

OpenStudy (paxpolaris):

\[\implies \frac 3 2 = h \times \frac94 \]

OpenStudy (paxpolaris):

\[h = \frac{ 3 }{ 2 } \div \frac{ 9 }{ 4 }\]

OpenStudy (paxpolaris):

\[h=\frac{ 3 }{ 2 }\times \frac{ 4 }{ 9 }\]

OpenStudy (paxpolaris):

\[h=\frac{ \cancel3 }{ \cancel2 }\times \frac{ \cancel4\ 2}{\cancel9\ 3}=\frac23\]

OpenStudy (paxpolaris):

the height will be 2/3 yards if we use all the blu paint